OpenStudy (anonymous):

A collection of dimes and quarters has a total value of five dollars and contains 29 coins. How much of each kind of coin are there in the collection? A. 15 dimes and 14 quarters B. 14 dimes and 15 quarters C. 10 dimes and 19 quarters D. 19 dimes and 10 quarters

OpenStudy (anonymous):

i think i answered a question just like this to you before q = amount of quarters d = amount of dimes .25q+.1d = 5 q+d = 29 solve the system of equations

OpenStudy (anonymous):

A 15 & 14 ?

OpenStudy (anonymous):

multiply the first equation by 20 and the second equation by 2 to solve by elimination easier 5q + 2d = 100 2q + 2d = 58 subtract the bottom one from the top 3q = 42 q = 14 plug your value for q back into q+d=29 14 +d = 19 d = 15 so 14 quarters and 15 dimes

OpenStudy (anonymous):

so yes, A is the answer

OpenStudy (anonymous):

thanks. (:

OpenStudy (anonymous):

you're welcome!